Ian is 6 years older than ryan.Six years ago he was twice as old as him.how old is each now?
Answer by Cromlix(4381)

Ian is 6 years older than Ryan.
I = R + 6
Six years ago he was twice as old as him.
I - 6 = 2(R - 6)
Replace I with R + 6 from top equation.
R + 6 - 6 = 2(R - 6)
R = 2R - 12
R - 2R = -12
-R = -12
R = 12
Currently:
Ryan is 12 years old
Ian is 18 years old.
Hope this helps.
:-)